# Explicit Second-Order Min-Max Optimization: Practical Algorithms and Complexity Analysis

### v5

@@ 1 Introduction
Our approach builds on existing algorithmic components: (i) the overall scheme is inspired by Lin and Jordan (2025); (ii) the subproblem reformulation is inspired by Adil et al. (2022); (iii) the use of the Newton method and the corresponding iteration complexity bound of $O(\log\log(1/\epsilon))$ extend the analysis of Conn et al. (2000); Cartis et al. (2011); (iv) the subsampling techniques are inspired by Xu et al. (2020). However, we remark that combining these components to yield a theoretically sound and practically efficient algorithm is far from straightforward.
While Adil et al. (2022) showed how to formulate the subproblem as a root-finding problem, a central challenge remains: how to solve such problems efficiently. Techniques such as the generalized conjugate gradient method with Lanczos processes—which are effective in the minimization setting—do not directly extend to the min-max setting. In this context, Huang et al. (2022) proposed an alternative formulation (see Eq. (16) in their paper) and applied a damped Newton method. However, they did not provide theoretical guarantees, even in a local sense, due to the absence of key properties such as the non-singularity of the Jacobian matrix. Even when adopting the root-finding formulation from Adil et al. (2022), identifying a suitable solver is highly nontrivial. At first glance, bisection might seem preferable due to its global convergence guarantee, unlike Newton’s method. One of our major contributions is to rigorously justify the use of the Newton method by extending the analysis of Conn et al. (2000) and Cartis et al. (2011) from the minimization setting to the min-max setting. Moreover, it remains unclear if the existing inexactness conditions from Xu et al. (2020) can be directly combined with our overall scheme while sacrificing neither theoretical iteration complexity nor practical efficiency. For example, the proof of Lemma 4.3 requires separate treatment of $|\Delta\textbf{z}_{k-1}|\geq 1$ and $|\Delta\textbf{z}_{k-1}|<1$, which was unnecessary in previous works either due to stronger assumptions as in Lin and Jordan (2025) or the simpler analysis appropriate for a minimization setting as in Xu et al. (2020).
@@ 4.3 Inexact subproblem solving and complexity analysis
We now clarify how to obtain an inexact solution of the nonlinear equation problem in Eq. (4.1) such that Condition 4.2 holds. In particular, we present a new subroutine that solves each subproblem using a single Schur decomposition and only $O(\log\log(1/\epsilon))$ calls to a linear system solver in a quasi-upper-triangular system. This gives a total complexity of $O((m+n)^{\omega}\epsilon^{-2/3}+(m+n)^{2}\epsilon^{-2/3}\log\log(1/\epsilon))$ which outperforms that of $O((m+n)^{\omega}\epsilon^{-2/3}\log\log(1/\epsilon))$ achieved by the best known line-search-based method.
For simplicity, we omit the subscript and rewrite the nonlinear equation problem in Eq. (4.1) as
$F(\hat{\mathbf{z}})+J(\hat{\mathbf{z}})\Delta\mathbf{z}+6\rho\|\Delta\mathbf{z}\|\Delta\mathbf{z}=\textbf{0}.$
This is then equivalent to finding a pair $(\Delta\mathbf{z},\lambda)$ for which
@@
and obtain from Eq. (4.9) and the above definition that the solution of $(\Delta\mathbf{z},\lambda)$ we are looking for depends upon the nonlinear equality $\lambda=6\rho\|\Delta\mathbf{z}(\lambda)\|$. For convenience, we define $\psi(\lambda)=\|\Delta\mathbf{z}(\lambda)\|^{2}$ and examine $\psi(\lambda)$ in the following proposition.
@@
If $F(\hat{\mathbf{z}})\neq\textbf{0}$, the function $\psi(\lambda)$ is strictly decreasing and convex when $\lambda>0$.
@@
$(J(\hat{\mathbf{z}})+\lambda I)\nabla_{\lambda}\Delta\mathbf{z}(\lambda)+\Delta\mathbf{z}(\lambda)=\textbf{0},\quad(J(\hat{\mathbf{z}})+\lambda I)\nabla^{2}_{\lambda\lambda}\Delta\mathbf{z}(\lambda)+2\nabla_{\lambda}\Delta\mathbf{z}(\lambda)=\textbf{0},$
@@
The nonlinear equality $\lambda=6\rho\|\Delta\mathbf{z}(\lambda)\|$ is equivalent to $\lambda=6\rho\sqrt{\psi(\lambda)}$ which can be reformulated as the following one-dimensional nonlinear equation problem:
@@
Suppose $F(\hat{\mathbf{z}})\neq\textbf{0}$. Then, we have the function $\phi(\lambda)$ is strictly decreasing and convex when $\lambda>0$. Its first-order derivative is
$\phi^{\prime}(\lambda)=-\tfrac{\Delta\mathbf{z}(\lambda)^{\top}(J(\hat{\mathbf{z}})+\lambda I)^{-1}\Delta\mathbf{z}(\lambda)}{\|\Delta\mathbf{z}(\lambda)\|}-\tfrac{1}{6\rho}.$
@@
$\phi^{\prime}(\lambda)=\tfrac{1}{2}\tfrac{\psi^{\prime}(\lambda)}{\sqrt{\psi(\lambda)}}-\tfrac{1}{6\rho},\quad\phi^{\prime\prime}(\lambda)=\tfrac{1}{4}\tfrac{2\psi^{\prime\prime}(\lambda)\psi(\lambda)-(\psi^{\prime}(\lambda))^{2}}{(\psi(\lambda))^{3/2}}.$
By Proposition 4.4, we have
$\phi^{\prime}(\lambda)=-\tfrac{\Delta\mathbf{z}(\lambda)^{\top}(J(\hat{\mathbf{z}})+\lambda I)^{-1}\Delta\mathbf{z}(\lambda)}{\|\Delta\mathbf{z}(\lambda)\|}-\tfrac{1}{6\rho}<0,\quad\textnormal{when }\lambda>0.$
We also have
$\phi^{\prime\prime}(\lambda)$ $=$ $\tfrac{\|\Delta\mathbf{z}(\lambda)\|^{2}(\|(J(\hat{\mathbf{z}})+\lambda I)^{-1}\Delta\mathbf{z}(\lambda)\|^{2}+2\Delta\mathbf{z}(\lambda)^{\top}(J(\hat{\mathbf{z}})+\lambda I)^{-2}\Delta\mathbf{z}(\lambda))-(\Delta\mathbf{z}(\lambda)^{\top}(J(\hat{\mathbf{z}})+\lambda I)^{-1}\Delta\mathbf{z}(\lambda))^{2}}{\|\Delta\mathbf{z}(\lambda)\|^{3}}$
$\geq$ $\tfrac{\|\Delta\mathbf{z}(\lambda)\|^{2}\|(J(\hat{\mathbf{z}})+\lambda I)^{-1}\Delta\mathbf{z}(\lambda)\|^{2}-(\Delta\mathbf{z}(\lambda)^{\top}(J(\hat{\mathbf{z}})+\lambda I)^{-1}\Delta\mathbf{z}(\lambda))^{2}}{\|\Delta\mathbf{z}(\lambda)\|^{3}}$
$\geq$ $0,\quad\textnormal{when }\lambda>0,$
@@
$J(\hat{\mathbf{z}})=QUQ^{-1}.$
A typical iteration of the Newton method for finding such root replaces the current iterate $\lambda^{j}>0$ with the improved estimate $\lambda^{j+1}$ for which
$\lambda^{j+1}=\lambda^{j}-\tfrac{\phi(\lambda^{j})}{\phi^{\prime}(\lambda^{j})}.$ (4.12)
@@
Suppose that the iterates $\{\lambda^{j}\}_{j\geq 0}$ are generated by the Newton method with $\lambda^{0}>0$ with $\phi(\lambda^{0})\geq 0$. Then, we have $\lambda^{j}>0$ and $\phi(\lambda^{j})\geq 0$. As a consequence, the iterates converge monotonically towards the unique solution $\lambda^{\star}$. The convergence rate is globally $Q$-linear with a factor at least $1-\phi^{\prime}(\lambda^{\star})/\phi^{\prime}(\lambda^{0})$ and is ultimately $Q$-quadratic.
Proof. Proposition 4.5 guarantees that $\phi^{\prime}(\lambda^{0})<0$ when $\lambda>0$, and it follows from Eq. (4.12) that $\lambda^{1}>\lambda^{0}>0$. Proposition 4.5 also guarantees the convexity and differentiability of $\phi$ which together with Eq. (4.12) implies that
$\phi(\lambda^{1})\geq\phi(\lambda^{0})+(\lambda^{1}-\lambda^{0})\phi^{\prime}(\lambda^{0})=0.$
Repeating this argument yields that $\lambda^{j}>0$ and $\phi(\lambda^{j})\geq 0$. In addition, $\phi$ is strictly decreasing. Thus, the iterates converge monotonically towards the unique solution.
Suppose that $\lambda^{\star}$ is the unique solution. Then, the mean value theorem implies
$\phi(\lambda^{j})=\phi(\lambda^{\star})+(\lambda^{j}-\lambda^{\star})\phi^{\prime}(\tilde{\lambda})=(\lambda^{j}-\lambda^{\star})\phi^{\prime}(\tilde{\lambda}),\quad\textnormal{for some }\tilde{\lambda}\in(\lambda^{j},\lambda^{\star}).$
Combining this with Eq. (4.12) yields
$|\lambda^{\star}-\lambda^{j+1}|=\left|\lambda^{\star}-\lambda^{j})\left(1-\tfrac{\phi^{\prime}(\tilde{\lambda})}{\phi^{\prime}(\lambda^{j})}\right)\right|\leq|\lambda^{\star}-\lambda^{j}|\cdot\left|1-\tfrac{\phi^{\prime}(\tilde{\lambda})}{\phi^{\prime}(\lambda^{j})}\right|.$
Proposition 4.5 guarantees that $\phi^{\prime}(\lambda)$ is increasing and strictly less than 0. Thus, we have $\phi^{\prime}(\lambda^{0})\leq\phi^{\prime}(\lambda^{j})\leq\phi^{\prime}(\tilde{\lambda})\leq\phi^{\prime}(\lambda^{\star})<0$ which implies that
$0\leq 1-\tfrac{\phi^{\prime}(\tilde{\lambda})}{\phi^{\prime}(\lambda^{j})}\leq 1-\tfrac{\phi^{\prime}(\lambda^{\star})}{\phi^{\prime}(\lambda^{0})}<1.$
Putting everything together then implies that the convergence rate is globally $Q$-linear with a factor at least $1-\phi^{\prime}(\lambda^{\star})/\phi^{\prime}(\lambda^{0})$. The asymptotic $Q$-quadratic convergence of the Newton iteration follows because the Jacobian of $\phi$ is nonsingular at the unique solution $\lambda^{\star}$ (i.e., $\phi^{\prime}(\lambda^{\star})\neq 0$). $\Box$
@@
Our reformulation closely follows Adil et al. (2022) who proposed a bisection method as a solver and proved a complexity bound of $O((m+n)^{\omega}\epsilon^{-2/3}+(m+n)^{2}\epsilon^{-2/3}\log(1/\epsilon))$ where $\omega\approx 2.3728$ is the matrix multiplication constant (Demmel et al., 2007). We also note that Huang et al. (2022) reformulated the same subproblem as a two-dimensional nonlinear equation (see Huang et al. (2022, Eq.(16))) and proposed a damped Newton method as a solver but without any complexity bound guarantee. Our finding is that the Newton method outperforms bisection and damped Newton in that it leverages the specific structure of Eq. (4.9) to yield not only an improved complexity bound but a practical implementation.
@@
Our analysis in Theorem 4.6 is related to Cartis et al. (2011, Theorem 6.3). The key difference lies in how we adapt their techniques to settings involving asymmetric Jacobian matrices. Specifically,  Cartis et al. (2011) analyze the function $\phi(\lambda)=1/\sqrt{\psi(\lambda)}-6\rho/\lambda$ and show that it is concave when the second-order information is symmetric. However, this function will not be concave in the presence of asymmetry. We address this issue by proposing a different function, $\phi(\lambda)=\sqrt{\psi(\lambda)}-\lambda/(6\rho)$, which we can prove remains concave even when the Jacobian matrix is asymmetric. This change is crucial because it allows us to retain the theoretical convergence guarantees of the Newton method without damping, even in the non-symmetric case.

### v6

@@ 1 Introduction
Our approach builds on existing algorithmic components: (i) the overall scheme is inspired by Lin and Jordan (2025); (ii) the subproblem reformulation is inspired by Adil et al. (2022); (iii) the use of the safeguarded Newton method and the corresponding iteration complexity bound of $O(\log\log(1/\epsilon))$ extend the analysis of Conn et al. (2000); Cartis et al. (2011); (iv) the subsampling techniques are inspired by Xu et al. (2020). However, we remark that combining these components to yield a theoretically sound and practically efficient algorithm is far from straightforward.
While Adil et al. (2022) showed how to formulate the subproblem as a root-finding problem, a central challenge remains: how to solve such problems efficiently. Techniques such as the generalized conjugate gradient method with Lanczos processes—which are effective in the minimization setting—do not directly extend to the min-max setting. In this context, Huang et al. (2022) proposed an alternative formulation (see Eq. (16) in their paper) and applied a damped Newton method. However, they did not provide theoretical guarantees, even in a local sense, due to the absence of key properties such as the non-singularity of the Jacobian matrix. Even when adopting the root-finding formulation from Adil et al. (2022), identifying a suitable solver is highly nontrivial. At first glance, bisection might seem preferable due to its global convergence guarantee, unlike Newton’s method. One of our major contributions is to rigorously justify the use of the safeguarded Newton method by extending the analysis of Conn et al. (2000) and Cartis et al. (2011) from the minimization setting to the min-max setting. Moreover, it remains unclear if the existing inexactness conditions from Xu et al. (2020) can be directly combined with our overall scheme while sacrificing neither theoretical iteration complexity nor practical efficiency. For example, the proof of Lemma 4.3 requires separate treatment of $|\Delta\textbf{z}_{k-1}|\geq 1$ and $|\Delta\textbf{z}_{k-1}|<1$, which was unnecessary in previous works either due to stronger assumptions as in Lin and Jordan (2025) or the simpler analysis appropriate for a minimization setting as in Xu et al. (2020).
@@ 4.3 Inexact subproblem solving and complexity analysis
[…]
For simplicity, we rewrite the nonlinear equation problem in Eq. (4.1) as
$F(\hat{\mathbf{z}})+J(\hat{\mathbf{z}})\Delta\mathbf{z}+6\rho\|\Delta\mathbf{z}\|\Delta\mathbf{z}=\mathbf{0}.$
This is equivalent to finding a pair $(\Delta\mathbf{z},\lambda)$ for which
@@
and obtain from Eq. (4.9) and the above definition that the solution $(\Delta\mathbf{z},\lambda)$ we are looking for depends upon the nonlinear equality $\lambda=6\rho\|\Delta\mathbf{z}(\lambda)\|$. For convenience, we define $\psi(\lambda)=\|\Delta\mathbf{z}(\lambda)\|^{2}$ and examine $\psi(\lambda)$ in the following proposition.
@@
If $F(\hat{\mathbf{z}})\neq\mathbf{0}$ and $v^{\top}J(\hat{\mathbf{z}})v\geq 0$ for all $v$, then $\psi(\lambda)$ is strictly decreasing on $(0,\infty)$.
@@
$(J(\hat{\mathbf{z}})+\lambda I)\nabla_{\lambda}\Delta\mathbf{z}(\lambda)+\Delta\mathbf{z}(\lambda)=\mathbf{0},\quad(J(\hat{\mathbf{z}})+\lambda I)\nabla^{2}_{\lambda\lambda}\Delta\mathbf{z}(\lambda)+2\nabla_{\lambda}\Delta\mathbf{z}(\lambda)=\mathbf{0}.$
@@
Putting these pieces together yields the desired expressions of $\psi^{\prime}(\lambda)$ and $\psi^{\prime\prime}(\lambda)$.
It remains to show that $\psi(\lambda)$ is strictly decreasing on $(0,\infty)$. For $0<\lambda_{1}<\lambda_{2}$, we have
@@
This implies
$J(\hat{\mathbf{z}})(\Delta\mathbf{z}(\lambda_{1})-\Delta\mathbf{z}(\lambda_{2}))+\lambda_{1}\Delta\mathbf{z}(\lambda_{1})-\lambda_{2}\Delta\mathbf{z}(\lambda_{2})=\mathbf{0}.$
Taking the inner product with $\Delta\mathbf{z}(\lambda_{1})-\Delta\mathbf{z}(\lambda_{2})$ yields
$(\Delta\mathbf{z}(\lambda_{1})-\Delta\mathbf{z}(\lambda_{2}))^{\top}J(\hat{\mathbf{z}})(\Delta\mathbf{z}(\lambda_{1})-\Delta\mathbf{z}(\lambda_{2}))+\lambda_{1}\|\Delta\mathbf{z}(\lambda_{1})\|^{2}-(\lambda_{1}+\lambda_{2})\Delta\mathbf{z}(\lambda_{1})^{\top}\Delta\mathbf{z}(\lambda_{2})+\lambda_{2}\|\Delta\mathbf{z}(\lambda_{2})\|^{2}=0.$
Since $v^{\top}J(\hat{\mathbf{z}})v\geq 0$ for all $v$, we have
$\lambda_{1}\|\Delta\mathbf{z}(\lambda_{1})\|^{2}-(\lambda_{1}+\lambda_{2})\Delta\mathbf{z}(\lambda_{1})^{\top}\Delta\mathbf{z}(\lambda_{2})+\lambda_{2}\|\Delta\mathbf{z}(\lambda_{2})\|^{2}\leq 0.$
This together with $\Delta\mathbf{z}(\lambda_{1})^{\top}\Delta\mathbf{z}(\lambda_{2})\leq\|\Delta\mathbf{z}(\lambda_{1})\|\|\Delta\mathbf{z}(\lambda_{2})\|$ and $\lambda_{2}>\lambda_{1}$ yields that $\|\Delta\mathbf{z}(\lambda_{1})\|>\|\Delta\mathbf{z}(\lambda_{2})\|$ whenever $F(\hat{\mathbf{z}})\neq\mathbf{0}$. This yields the desired result. $\Box$
The nonlinear equality $\lambda=6\rho\|\Delta\mathbf{z}(\lambda)\|$ is equivalent to $\lambda=6\rho\sqrt{\psi(\lambda)}$, which can be reformulated as the following one-dimensional nonlinear equation problem:
@@
Suppose that $F(\hat{\mathbf{z}})\neq\mathbf{0}$ and $v^{\top}J(\hat{\mathbf{z}})v\geq 0$ for all $v$. Then, $\phi(\lambda)$ is strictly decreasing on $(0,\infty)$. Its first-order derivative is
$\phi^{\prime}(\lambda)=-\tfrac{\Delta\mathbf{z}(\lambda)^{\top}(J(\hat{\mathbf{z}})+\lambda I)^{-1}\Delta\mathbf{z}(\lambda)}{\|\Delta\mathbf{z}(\lambda)\|}-\tfrac{1}{6\rho}<0,\quad\textnormal{for all }\lambda>0.$
Consequently, Eq. (4.11) admits a unique solution $\lambda^{\star}>0$.
@@
$\phi^{\prime}(\lambda)=\tfrac{1}{2}\tfrac{\psi^{\prime}(\lambda)}{\sqrt{\psi(\lambda)}}-\tfrac{1}{6\rho}.$
By Proposition 4.4, $\psi^{\prime}(\lambda)<0$ for all $\lambda>0$. This implies $\phi^{\prime}(\lambda)<0$. Since $\phi$ is continuous and strictly decreasing, the root is unique. The existence of such root follows from $\phi(\lambda)\to$-$\infty$ as $\lambda\to+\infty$ and the fact that $\phi(\lambda)>0$ if $\lambda$ is sufficiently close to 0. $\Box$
The required solution is the unique positive root to Eq. (4.11). Let $\lambda^{0}>0$ be given with $\phi(\lambda^{0})>0$, we first perform one Schur decomposition: $J(\hat{\mathbf{z}})=QUQ^{-1}$. Then, a typical iteration of the Newton method for finding such root replaces the current iterate $\lambda^{j}>0$ with the candidate $\tilde{\lambda}^{j+1}$ for which
$\tilde{\lambda}^{j+1}=\lambda^{j}-\tfrac{\phi(\lambda^{j})}{\phi^{\prime}(\lambda^{j})},$ (4.12)
[…]
In terms of convergence, the Newton method is not globally convergent for the general asymmetric case. We can adopt a safeguarded Newton-bisection scheme: given an interval $L^{j},U^{j}]$ satisfying $\phi(L^{j})\geq 0\geq\phi(U^{j})$ and a current iterate $\lambda^{j}\in[L^{j},U^{j}]$, we compute $\tilde{\lambda}^{j+1}$ using (4.12). If $\tilde{\lambda}^{j+1}\in(L^{j},U^{j})$, we accept it and set $\lambda^{j+1}=\tilde{\lambda}^{j+1}$; otherwise we take a bisection step and set $\lambda^{j+1}=(L^{j}+U^{j})/2$. We then update the bracket using the sign of $\phi(\lambda^{j+1})$.
Theorem 4.6
Suppose that $F(\hat{\mathbf{z}})\neq\mathbf{0}$ and $v^{\top}J(\hat{\mathbf{z}})v\geq 0$ for all $v$. Let $\lambda^{\star}>0$ be the unique solution to Eq. (4.11). Suppose that the safeguarded Newton-bisection method is initialized with $L^{0},U^{0}]$ such that $\phi(L^{0})\geq 0\geq\phi(U^{0})$. Then, $\lambda^{j}\to\lambda^{\star}$ globally. There also exists a constant $r>0$ such that:
  1. 1.
    After at most $O(\log((U^{0}-L^{0})/r))$ iterations, the iterates enter the local Newton region $|\lambda-\lambda^{\star}|\leq r$.
  2. 2.
    Once inside this region, all Newton steps are accepted and the convergence is $Q$-quadratic, requiring $O(\log\log(1/\epsilon))$ additional iterations to achieve $|\lambda^{j}-\lambda^{\star}|\leq\epsilon$.
The total number of iterations is $O(\log(1/r))+O(\log\log(1/\epsilon))$ where $r>0$ is independent of $\epsilon$.
Proof. By Proposition 4.5, $\phi$ is continuous and strictly decreasing. Thus, the bracketing interval remains valid throughout the iterations and contains $\lambda^{\star}$. Whenever the iterate $\tilde{\lambda}^{j+1}$ leaves the bracket, the method takes a bisection step, which halves the bracket length. After at most $O(\log((U^{0}-L^{0})/r))$ safeguarding steps, the bracket width is at most $r$, and the iterate satisfies $|\lambda^{j}-\lambda^{\star}|\leq r$.
By Proposition 4.5, we have $\phi(\lambda^{\star})>0$ and $\phi^{\prime}$ is Lipschitz around the unique root $\lambda^{\star}$. Standard one-dimensional Newton analysis implies that there exists $r>0$ such that if $|\lambda^{j}-\lambda^{\star}|\leq r$, the Newton step stays within the neighborhood, is accepted, and the iterates converge $Q$-quadratically. The number of Newton iterations required to reach the tolerance $\epsilon$ is thus $O(\log\log(1/\epsilon))$. Combining the safeguarding phase and the local Newton phase yields the desired result. $\Box$
@@
We let $\eta=\lambda^{2}$ and define
$\Psi(\eta)\triangleq F(\hat{\mathbf{z}})^{\top}(\eta I+S)^{-1}F(\hat{\mathbf{z}}),\quad\Phi(\eta)\triangleq\Psi(\eta)-\tfrac{\eta}{36\rho^{2}}.$
Then, solving $\phi(\lambda)=0$ is equivalent to solving $\Phi(\eta)=0$ with $\lambda^{\star}=\sqrt{\eta^{\star}}$.
Proposition 4.7
@@
Proof. Since $S\succeq 0$, the map $\eta\mapsto(\eta I+S)^{-1}$ is operator monotone decreasing and operator convex on $(0,\infty)$, which implies that $\Psi(\eta)$ is strictly decreasing and convex. Subtracting the linear term $\eta/(36\rho^{2})$ preserves convexity and strict monotonicity. This yields the desired result. $\Box$
@@
[…]
$\eta^{j+1}=\eta^{j}-\tfrac{\Phi(\eta^{j})}{\Phi^{\prime}(\eta^{j})},\quad\Phi(\eta^{0})\geq 0.$
Then, $\eta^{j}$ increases monotonically and converges to the unique solution $\eta^{\star}$. Moreover, the rate is globally $Q$-linear with a factor at least $1-\Phi^{\prime}(\lambda^{\star})/\Phi^{\prime}(\lambda^{0})$ and is ultimately $Q$-quadratic. Equivalently, we have $\lambda^{j}=\sqrt{\eta^{j}}$ converges to $\lambda^{\star}=\sqrt{\eta^{\star}}$, and the number of iterations to achieve $|\lambda^{j}-\lambda^{\star}|\leq\epsilon$ is $O(\log\log(1/\epsilon))$.
Proof. By Proposition 4.7, we have $\Phi^{\prime}(\eta^{0})<0$, and it follows from the update formula that $\eta^{1}>\eta^{0}>0$. Proposition 4.7 also guarantees the convexity and differentiability of $\phi$ which together with the update formula implies that
$\Phi(\eta^{1})\geq\phi(\eta^{0})+(\eta^{1}-\eta^{0})\Phi^{\prime}(\eta^{0})=0.$
Repeating this argument yields that $\eta^{j}>0$ and $\Phi(\eta^{j})\geq 0$. In addition, $\Phi$ is strictly decreasing. Thus, the iterates converge monotonically towards the unique solution. Suppose that $\eta^{\star}$ is the unique solution. Then, the mean value theorem implies
$\Phi(\eta^{j})=\Phi(\eta^{\star})+(\eta^{j}-\eta^{\star})\Phi^{\prime}(\tilde{\eta})=(\eta^{j}-\eta^{\star})\Phi^{\prime}(\tilde{\eta}),\quad\textnormal{for some }\tilde{\eta}\in(\eta^{j},\eta^{\star}).$
Combining this with the update formula yields
$|\eta^{\star}-\eta^{j+1}|=\left|(\eta^{\star}-\eta^{j})\left(1-\tfrac{\Phi^{\prime}(\tilde{\eta})}{\Phi^{\prime}(\eta^{j})}\right)\right|\leq|\eta^{\star}-\eta^{j}|\cdot\left|1-\tfrac{\Phi^{\prime}(\tilde{\eta})}{\Phi^{\prime}(\eta^{j})}\right|.$
Since $\Phi$ is convex, we have $\Phi^{\prime}(\lambda^{0})\leq\Phi^{\prime}(\lambda^{j})\leq\Phi^{\prime}(\tilde{\lambda})\leq\Phi^{\prime}(\lambda^{\star})<0$ which implies that
$0\leq 1-\tfrac{\Phi^{\prime}(\tilde{\eta})}{\Phi^{\prime}(\eta^{j})}\leq 1-\tfrac{\Phi^{\prime}(\eta^{\star})}{\Phi^{\prime}(\eta^{0})}<1.$
Putting everything together then implies that the convergence rate is globally $Q$-linear with a factor at least $1-\Phi^{\prime}(\eta^{\star})/\Phi^{\prime}(\eta^{0})$. The asymptotic $Q$-quadratic convergence of the Newton iteration follows. This together with $\lambda=\sqrt{\eta}$ yields the desired result. $\Box$
@@
Our reformulation closely follows Adil et al. (2022) who proposed a bisection method as a solver and proved a complexity bound of $O((m+n)^{\omega}\epsilon^{-2/3}+(m+n)^{2}\epsilon^{-2/3}\log(1/\epsilon))$ where $\omega\approx 2.3728$ is the matrix multiplication constant (Demmel et al., 2007). Our new schemes have achieved the better worst-case dependence on $\epsilon$ thanks to a significantly faster quadratic convergence phase.
@@
Our analysis is related to Cartis et al. (2011, Theorem 6.3). The key difference lies in how asymmetry affects curvature properties of one-dimensional reformulations. In general asymmetric settings, neither concavity nor convexity of $\phi(\lambda)$ is guaranteed, and thus the Newton method requires safeguarding for global convergence. In contrast, under the antisymmetry condition, $\eta=\lambda^{2}$ restores convexity of the scalar equation and yields a globally convergent unmodified Newton scheme.